# Direct numerical simulations of a cylinder cutting a vortex

### v1

@@ Abstract
The interaction between a vortex and an impacting body which is oriented normally to it is complex due to the interaction of inviscid and viscous three-dimensional mechanisms. To model this process, we conducted direct numerical simulations of a thin cylinder intersecting a columnar vortex oriented normally to it. We identified the various stages of the interaction, including the separation of the boundary layer and the generation of secondary vorticity that interacts with the primary vortex. By varying the impact parameter and the Reynolds number, we were able to distinguish the two regimes of interaction mentioned in the literature: the weak and strong vortex regimes. We found that low impact parameters, representing strong vortices, led to ejection and interaction of secondary vorticity from the cylinder’s boundary layer, while high impact parameters, representing weak vortices, led to approximately inviscid interaction of the cylinder with the primary vortex through deformations. We did not find a significant effect of the Reynolds number in the overall phenomenology, even if larger Reynolds numbers lead to the formation of increasingly smaller and more intense vortex structures in the parameter range studied. Finally, we analyzed the hydrodynamic force curves on the cylinder, showing that intense forces could be locally generated for some parameter regimes, but that the average force on the cylinder did not deviate substantially from baseline cases where no vortex was present. Our results shed light on the underlying mechanisms of vortex-body interactions and their dependence on various parameters.
@@ I Introduction
BVI holds widespread significance in various engineering applications, spanning wind turbines, helicopter rotors, marine propellers, and industrial fans [6], and due to its practical applications has been the workhorse of research in vortex line cutting in fluid dynamics. The majority of research on BVI concentrates on aerodynamic blades intersecting a straight vortex tube. This specific BVI scenario, referred to as perpendicular or normal BVI, occurs when the vortex and blade axes are perpendicular, and the blade cleaves through the vortex tube. Normal BVI is commonly encountered in applications such as helicopter rotors and wind turbines, where rotating blades confront vortices originating from prior blades or generated by their own motion [9, 10, 11, 12, 13]. We note that other orientations of the blade in relation to the vortex are possible, resulting in other types of BVI: grazing BVI, where the blade is perpendicular to the tube, but it moves in the direction of the core, and parallel BVI [14], where the blade is parallel to the tube, and cuts the tube throughout its extent rather than at a single point. These are outside the scope of our study which focuses on swift cutting of the core.
@@
Depending on both the response of the primary vortex to the body and about the interaction of the vortex with the body’s boundary layer, response regimes can be constructed depending on the value the dimensionless parameters take. The response regimes are commonly referred to as the weak vortex regime, the strong vortex regime, and the bending regime. The weak vortex regime occurs for thin bodies $(T=D/\sigma_{0}\lesssim 1)$ when the vortex is sufficiently weak $(I_{P}>0.2)$. This regime is characterized with very minimal boundary layer separation and ejection until after the body has penetrated the primary vortex [22]. Once the boundary layer has separated, the separated vorticity becomes entrained within the primary vortex and does not wrap around it, instead spreading away from the body and into the vortex core.
@@
Inviscid filament models are the oldest method used to simulate BVI, due to their straightforwardness and effectiveness [24, 23, 3]. Essentially, filament models assume that vorticity is concentrated within a tiny filament, and use the Biot-Savart line integral to calculate the resulting velocity field. However, when the integral is evaluated on the filament itself, the result is singular and cannot be used to determine the self-induced velocity of the vortex. To overcome this issue, the domain of integration is typically limited to a specific range that varies with the size and vorticity field of the vortex [24]. Another way to study vortex-body interaction is to use Helmholtz’s law [25, 26]. This approach has been extended to account for variations in the vortex core area during wave propagation. However, a no-penetration condition on the body surface must be ensured, which can be achieved using a vortex sheet [3]. Filament models are particularly useful for predicting how the body will penetrate the vortex or how the vortex will react to the motion of the body [23, 27, 3]. These models have also been used to estimate the pressure on the body surface during the interaction with a vortex, but only up to a certain distance from the body [27]. The distance may vary depending on the impact parameter.
While filament models have yielded promising results, they are limited by the inviscid assumption, and cannot be used to analyze boundary layer separation induced by the vortex. Understanding this phenomena is crucial, because when the body’s boundary layer separates, the ejected vorticity from the boundary layer vorticity interacts with the primary vortex to form secondary vortices, which modifies the interaction process and causes different flow phenomenology in the weak and strong vortex regimes. Even if this process is highly three-dimensional, a number of studies which combine in two-dimensions have been conducted which have hinted at the type of complex flow phenomena which arise. Doligalski et al[28] analyzed a rectilinear vortex above a flat plate and found evidence of separation and ejection of fluid within the boundary layer into the external flow. Luton et al. [29] found that a weaker vortex or secondary vorticity can redistribute the vorticity of the boundary layer, leading to ejection of boundary layer fluid. A number of three-dimensional studies which combine filaments and viscous boundary layers have been conducted to study the detailed separation process and its evolution in the interaction [30, 31]. These show indications of singular behaviour when using vortex filaments, and cannot fully resolve the dynamics [31].
@@
As can be gathered from above, a detailed numerical study of the separation process in the weak vortex regime and the transition from weak to strong vortex regimes is missing. The use of DNS in vortex-body interaction studies has the potential to clarify previously unclear areas such as the viscous interaction of the body with the vortex [32]. Simulations also allow careful control of the Reynolds number and can assess questions of Reynolds number independence unanswered in experiments [17]. In this manuscript, we set to address these knowledge gaps by conducting three-dimensional direct numerical simulations (DNS) of tube cutting using state-of-the-art numerical techniques. We will simulate a simplified problem: a thin wire (cylinder) impacting normally a high-Reynolds number vortex, and we will study the interaction process by varying two parameters: the impact parameter $I_{P}$ and the circulation Reynolds number $Re_{\Gamma}$, to access the different regimes, and to measure the effect of the Reynolds number on the flow, and on the boundary layer vorticity in particular. By simulating a cylinder, we will enhance the generation of wake turbulence, and can assess how our results differ from those simulations which use blades.
We will focus on key phenomena observed during the interaction. These included the separation and ejection of the boundary layer from the body, the extent of bending in the primary vortex, the emergence from and impact of secondary vortices on the primary vortex, and the topology of major flow structures affecting the induced force on the cylinder, including possible reconnection phenomena. To analyze the flow field, we will employ vorticity visualizations, and calculate the induced force on the cylinder as it cuts through the vortex. We will identify the major flow structures influencing the force curve, and establish qualitative relationships between changes in the impact parameter and circulation Reynolds number.
@@ II Numerical methods
where u is the velocity, $t$ is time, $\rho$ is the fluid density, $\nu$ is the fluid kinematic viscosity and f is a body force originating from the the immersed boundary method (IBM) forcing used to model the cylinder. A schematic of the system is presented in Figure 1, which shows a cylinder of diameter $D$, with an axis parallel to the $z$-axis travels in the $x$ direction at a velocity $V$ towards a vortex of core-size $\sigma_{0}$ which has an axis parallel to the $y$ direction. This vortex is initialized at the initial time with a Lamb-Ossen (Gaussian) velocity profile with circulation $\Gamma$. We use the vortex circulation $\Gamma$ and the vortex radius $\sigma_{0}$ to non-dimensionalize the system.
@@
We fix the diameter of the cylinder as $0.95\sigma_{0}$, keeping the thickness parameter $T=D/\sigma_{0}\approx 1$ in the thin regime. We vary the impact parameter between $I_{P}\in(0.05,0.25)$. With our value of $T$, the degree of (inviscid) vortex bending that occurs before the ejection of boundary layer vorticity is determined by the impact parameter. The range of values of $I_{P}$ chosen, which allows us to explore the weak- and strong interaction regimes for low and high values of $I_{P}$ respectively. Finally, we also vary the circulation-based Reynolds numbers in the range $Re_{\Gamma}\in(1000,3000)$ to study the effect of boundary layer turbulence on the system dynamics. The associated cylinder Reynolds number $Re_{c}=VD/\nu$ is calculated as $Re_{c}=\Gamma I_{P}T/(2\pi\nu)=Re_{\Gamma}I_{P}T/(2\pi)$, and varies in the range $Re_{c}\in(7.96,119)$. Table 1 summarizes the control parameters used for all the runs in this manuscript. We note that while the values of $Re_{c}$ used here appear substantially lower than those used in Ref. [32], the reference uses a chord-based Reynolds number, while we use something analogous to a “thickness”-based Reynolds number leading to an order of magnitude discrepancy when comparing $Re_{c}$ to the blade Reynolds number in Ref. [32].
@@
Simulations are ran using the open-source code AFiD [33]. This code uses energy-conserving second order centered finite differences to spatially discretize the domain. Time marching is done through a third-order Runge-Kutta for the non-linear terms and a second order Crank-Nicolson for the viscous terms. The spatial resolution is taken as $624\times 624\times 1872$, meaning the grid is equally spaced in all directions, with a grid length $\sim 0.05\sigma_{0}\approx 0.05D$. Resolution adequacy is checked by monitoring numerical dissipation. The cylinder is incorporated through the immersed boundary method (IBM) using a moving-least-squares formulation [34]. The surface is discretized using $388800$ triangles. The triangle skewness is $<0.3$, and the average edge length is about $70\%$ of the grid spacing. This ensures adequate resolution without consuming excessive computational resources [34].
@@ III Flow Dynamics for Tube-Wire interaction
III.1 The weak and strong vortex regimes
We initially examine cases at $Re_{\Gamma}=1000$, varying $I_{P}$ to illustrate the difference between the strong and weak vortex regimes. Our analysis begins with visualizing $Q=\partial_{i}u_{j}\partial_{j}u_{i}$ [35], which enables us to focus on the primary vortex. A subsequent discussion of vorticity magnitude will emphasize the wire’s boundary layer and secondary vortices.
@@ III.1 The weak and strong vortex regimes
Figure 2: Visualization of the time evolution of $Q$ for a simulation in the strong vortex regime with $Re_{\Gamma}=1000$, $I_{P}=0.05$. Regions in blue indicate regions of strongly negative $Q$. Snapshots are separated $t\Gamma/D^{2}=200$ time-units.
Figure 3 depicts a similar volume rendering of $Q$ in the weak vortex regime ($Re_{\Gamma}=1000$, $I_{P}=0.25$). In this regime, the primary vortex does not “pre-sense” the wire, instead interacting with it solely when it is in close proximity to the vortex core. The vortex swiftly traverses the tube, causing significant deformation to the core. This deformation generates upwards and downwards-propagating waves. As time progresses, the wire completes its passage through the primary vortex tube, leaving behind a minor wake and a distorted tube. Surprisingly, despite the distortion, the “weak” vortex tube displays better resilience compared to the scenario involving a strong vortex. By the end of the BVI process, the waves have transversed the tube and a distorted vortex tube remains behind.
This means that unintuitively, at $Re_{\Gamma}=1000$, a vortex tube endures the “cut” of its core more effectively when encountering a weaker cutting object. This observation warrants further analysis, which we delve into shortly. Primarily, this outcome emerges because the interaction with a slower wire is prolonged, allowing the vortex more time to engage with secondary vorticity originating from the wire. Consequently, this extended interaction disrupts the primary vortex to a greater extent.
Figure 3: Visualization of the time evolution of $Q$ for a simulation in the weak vortex regime with $Re_{\Gamma}=1000$, $I_{P}=0.25$. Regions in blue indicate regions of strongly negative $Q$. Snapshots are separated $t\Gamma/D^{2}=40$ time-units.
@@
Figure 4: Visualization of $Q$ after the wire has passed through the cylinder axis for $Re_{\Gamma}=1000$, and $I_{P}=0.05$, $0.1$, $0.15$ and $0.25$ (left to right). Regions in blue indicate regions of strongly negative $Q$.
@@
To investigate the role of secondary vorticity in the vortex-wire interaction, our focus turns to vorticity visualization. Beginning with the analysis of the strong vortex regime at $I_{P}=0.05$, Figure 5 presents volumetric representations of the vorticity magnitude $|\omega|$ at $Re_{\Gamma}=1000$. These visuals align with the experimental images and descriptions in Ref. [17]: as the vortex approaches the tube, vorticity from its boundary layer detaches, and begins to curl around the primary tube. Even before the wire approaches the vortex core, this detached vorticity loops once around the main vortex, substantially distorting the core. A compression or stretching of the core occurs around the impact zone due to both the wire and the secondary vorticity, forming a neck region with heightened vorticity magnitude to conserve circulation. Subsequently, the detached vorticity propagates further along the vortex axis, nearly reaching the computational domain’s periodic boundaries. Upon completion of the wire’s cut through the vortex, the primary vortex remains significantly disrupted by the detached secondary vorticity, leaving behind a weakened core. A turbulent cloud of small scale vortices can also be observed when visualizing $|\omega|$. These are rapidly dissipated by viscosity.
@@
Figure 5: Visualization of the time evolution of $|\omega|$ for a simulation in the strong vortex regime with $Re_{\Gamma}=1000$, $I_{P}=0.05$. Snapshots are at $t\Gamma/D^{2}=740$, $1140$, $1540$ and $1740$.
@@
While these observed flow phenomena mirror those documented experimentally in Ref. [17], our findings expand upon the experimental groundwork in several crucial ways. Firstly, our simulations portray vorticity three-dimensionally rather than photographs of dye visualization. This representation enables us to illustrate the propagation of secondary vorticity through the axis, a phenomenon which is not clear in the experiments, as well as showing the new vortex filaments produced. Secondly, we demonstrate that secondary vorticity arises without the need to introduce turbulence sources either through background flow or vibrations of the structure. Even starting from simple initial conditions, such as a Lamb-Oseen vortex, the secondary vorticity originating solely from the wire’s boundary layer is enough to faithfully reproduce the significant disruptions in the primary vortex seen experimentally. Lastly, we note that our results are symmetrical around the cutting plane in our results due to the absence of axial flow. This shows that the presence of a weak axial flow does not change the phenomenology of strong vortex BVI very much.
In Figure 8, we provide further qualitative data with the vorticity magnitude plotted at the central $x$-$z$ plane for four time instants: before, during, and after core cutting. The first and second snapshots reveal significant detachment and wrapping of vorticity from the wire’s boundary layer before the wire reaches the core, leading to core deformation and vorticity profile redistribution. The third panel, just before impact, displays vorticity amplification due to tube stretching, coinciding with the interaction following a substantial core displacement from its original position. Finally, the fourth panel depicts the remnants of the vortex, significantly weakened and displaced six to seven vortex radii from its initial position. The final core circulation has been reduced to a quarter of its original strength.
@@
Figure 6: Visualization of the time evolution of $|\omega|$ for a simulation in the weak vortex regime with $Re_{\Gamma}=1000$, $I_{P}=0.05$. Snapshots are at $t\Gamma/D^{2}=540$, $940$, $1340$ and $2340$.

### v2

@@ Abstract
The interaction between a vortex and an impacting body which is oriented normally to it is complex due to the interaction of inviscid and viscous three-dimensional mechanisms. To model this process, direct numerical simulations of a thin cylinder intersecting a columnar vortex oriented normally to it are conducted. By varying the impact parameter and the Reynolds number, the two regimes of interaction mentioned in the literature are distinguished: the weak and strong vortex regimes. Low impact parameters, representing strong vortices, led to ejection and interaction of secondary vorticity from the cylinder’s boundary layer, while high impact parameters, representing weak vortices, led to approximately inviscid interaction of the cylinder with the primary vortex through deformations. No significant effect of the Reynolds number in the overall phenomenology is found, even if larger Reynolds numbers lead to the formation of increasingly smaller and more intense vortex structures in the parameter range studied. Finally, the hydrodynamic force curves on the cylinder are analyzed, showing that intense forces could be locally generated for some parameter regimes, but that the average force on the cylinder does not substantially deviate from baseline cases where no vortex was present. Our results shed light on the underlying mechanisms of vortex-body interactions and their dependence on various parameters.
@@ I Introduction
BVI holds widespread significance in various engineering applications, spanning wind turbines, helicopter rotors, marine propellers, and industrial fans [6], and due to its practical applications has been the workhorse of research in vortex line cutting in fluid dynamics. The majority of research on BVI concentrates on aerodynamic blades intersecting a straight vortex tube. This specific BVI scenario, referred to as perpendicular or normal BVI, occurs when the vortex and blade axes are perpendicular, and the blade cleaves through the vortex tube. Normal BVI is commonly encountered in applications such as helicopter rotors and wind turbines, where rotating blades confront vortices originating from prior blades or generated by their own motion [9, 10, 11, 12, 13]. Other orientations of the blade in relation to the vortex are possible, resulting in other types of BVI: grazing BVI, where the blade is perpendicular to the tube, but it moves in the direction of the core, and parallel BVI [14], where the blade is parallel to the tube, and cuts the tube throughout its extent rather than at a single point. These are outside the scope of our study which focuses on swift cutting of the core.
@@
Depending on both the response of the primary vortex to the body and about the interaction of the vortex with the body’s boundary layer, response regimes can be constructed depending on the value the dimensionless parameters take. The response regimes are commonly referred to as the weak vortex regime, the strong vortex regime, and the bending regime. In this manuscript, we will simulate cases in the weak and strong vortex regime, while the bending regime will not be investigated. However, we discuss here all three regimes for completeness.
The weak vortex regime occurs for thin bodies $(T=D/\sigma_{0}\lesssim 1)$ when the vortex is sufficiently weak $(I_{P}>0.2)$. This regime is characterized with very minimal boundary layer separation and ejection until after the body has penetrated the primary vortex [22]. Once the boundary layer has separated, the separated vorticity becomes entrained within the primary vortex and does not wrap around it, instead spreading away from the body and into the vortex core.
@@
Inviscid filament models are the oldest method used to simulate BVI, due to their straightforwardness and effectiveness but present well-known limitations: they require extensions to account for variations in the vortex core area and to ensure no-penetration conditions on the surfaces; and they are inviscid, which means that they cannot be used to analyze vortex-induced boundary layer separation [24, 25, 26, 23, 3], a crucial phenomena in the weak vortex regime. Nonetheless, they are useful for predicting how the body will penetrate the vortex or how the vortex will react to the motion of the body, and can also been used to estimate the pressure on the body surface during the interaction with a vortex up to a certain distance from the body [23, 27]. A number of extensions to the filament model which incorporate two-dimensional boundary layer dynamics have been conducted, which hint at the type of complex flow phenomena which arise [28, 29]. These were later extended to include three-dimensional boundary layers [30, 31]. However, they remain limited by the fact that the underlying filament models present singularities and cannot fully resolve the dynamics [31].
@@
As can be gathered from above, a detailed numerical study of the separation process in the weak vortex regime and the transition from weak to strong vortex regimes is missing. The use of DNS in vortex-body interaction studies has the potential to clarify previously unclear areas such as the viscous interaction of the body with the vortex [32]. Simulations also allow careful control of the Reynolds number and can assess questions of Reynolds number independence unanswered in experiments [17]. In this manuscript, we set to address these knowledge gaps by conducting three-dimensional direct numerical simulations (DNS) of tube cutting using state-of-the-art numerical techniques. We will simulate a simplified problem: a thin wire (cylinder) impacting normally a high-Reynolds number vortex. We will study the interaction process by varying two parameters: the impact parameter $I_{P}$ and the circulation Reynolds number $Re_{\Gamma}$. Values of $I_{P}$ will be chosen such that simulations can access the weak and strong vortex regimes, while varying $Re_{\Gamma}$ will modify the secondary vorticity of the cylinder, as well potentially triggering more hydrodynamical instabilities in the primary tubes which set in at values of $Re_{\Gamma}\sim\mathcal{O}(10^{3})$ [33]. We will focus on key flow phenomena observed during the interaction, calculate the induced force on the cylinder as it cuts through the vortex, and relate how major flow structures influence the force curve.
The manuscript is organized as follows: Section II details the numerical methods used in the manuscript. Section III presents the main results from the flow field, including the transition between strong and weak vortex regimes (III.1), and in-depth analysis of the strong (III.2) and weak vortex regimes (III.3). Section IV examines the force in the wire. Finally, Section V presents a summary of the findings as well as an outlook for future work.
@@ II Numerical methods
where u is the velocity, $t$ is time, $\rho$ is the fluid density, $\nu$ is the fluid kinematic viscosity and f is a body force originating from the the immersed boundary method (IBM) forcing used to model the cylinder. A schematic of the system is presented in Figure 1, which shows a cylinder of diameter $D$, with an axis parallel to the $z$-axis travels in the $x$ direction at a velocity $V$ towards a vortex of core-size $\sigma_{0}$ which has an axis parallel to the $y$ direction. This vortex is initialized at the initial time with a Lamb-Ossen (Gaussian) velocity profile with circulation $\Gamma$. The vortex circulation $\Gamma$ and the vortex radius $\sigma_{0}$ are used to non-dimensionalize the system.
@@
The diameter of the cylinder is fixed as $0.95\sigma_{0}$, keeping the thickness parameter $T=D/\sigma_{0}\approx 1$ in the thin regime. The impact parameter is varied between $I_{P}\in(0.05,0.25)$. With this value of $T$, the degree of (inviscid) vortex bending that occurs before the ejection of boundary layer vorticity is determined by the impact parameter. The range of values of $I_{P}$ chosen, which allows us to explore the weak- and strong interaction regimes for low and high values of $I_{P}$ respectively. Finally, the circulation-based Reynolds numbers are varied in the range $Re_{\Gamma}\in(1000,3000)$. This will impact the flow in two ways: the boundary layer of the cylinder will be thinner with higher vorticity, and the increased Reynolds number may also trigger hydrodynamical instabilities in the primary tube, as many instabilities in tubes set in at values of $Re_{\Gamma}$ which are between $1000$ and $3000$ [33].
The associated cylinder Reynolds number $Re_{c}=VD/\nu$ is calculated as $Re_{c}=\Gamma I_{P}T/(2\pi\nu)=Re_{\Gamma}I_{P}T/(2\pi)$, and varies in the range $Re_{c}\in(7.96,119)$. Table 1 summarizes the control parameters used for all the runs in this manuscript. We note that while the values of $Re_{c}$ used here appear substantially lower than those used in Ref. [32], the reference uses a chord-based Reynolds number, while this manuscript uses something analogous to a “thickness”-based Reynolds number leading to an order of magnitude discrepancy when comparing $Re_{c}$ to the blade Reynolds number in Ref. [32].
@@
Simulations are ran using the open-source code AFiD [34]. This code uses energy-conserving second order centered finite differences to spatially discretize the domain. Time marching is done through a third-order Runge-Kutta for the non-linear terms and a second order Crank-Nicolson for the viscous terms. The spatial resolution is taken as $624\times 624\times 1872$, meaning the grid is equally spaced in all directions, with a grid length $\sim 0.05\sigma_{0}\approx 0.05D$. Resolution adequacy is checked by monitoring numerical dissipation to be less than $2\%$. To increase the confidence in the mesh adequacy, Appendix A shows a comparison between a coarser mesh, the mesh used in this manuscript and a finer mesh for the $I_{P}=0.25$, $Re_{\Gamma}=3000$ case.
The cylinder is incorporated through the immersed boundary method (IBM) using a moving-least-squares formulation [35]. This algorithm has been used previously for turbulent simulations with rigid, moving and deforming boundaries, and has been well validated [36, 37, 38]. The surface is discretized using $388800$ triangles. The triangle skewness is $<0.3$, and the average edge length is about $70\%$ of the grid spacing. This ensures adequate resolution without consuming excessive computational resources [35].
@@ III Flow Dynamics for Tube-Wire interaction
III.1 The distinction between weak and strong vortex regimes
This subsection examines cases at $Re_{\Gamma}=1000$, varying $I_{P}$ to illustrate the difference between the strong and weak vortex regimes. The $Q$-criterium [39, 40] will be used to highlight the primary vortex. This is defined as:
$Q=-\frac{\partial u_{j}}{\partial x_{i}}\frac{\partial u_{i}}{\partial x_{j}}=\frac{1}{2}\omega^{2}-\sigma^{2}$ (3)
where $\omega^{2}=(\nabla\times\textbf{u})^{2}$ is the vorticity magnitude squared, and $\sigma_{ij}=\frac{1}{2}(\partial_{i}u_{j}+\partial_{j}u_{i})$ the strain tensor, with $\sigma^{2}=\sigma_{ij}\sigma_{ij}$. $Q$ represents the source term in the pressure-Poisson equation [41] and is useful for distinguishing rotation-dominated regions of the flow, which will tend to have positive values of $Q$, from strain-dominated regions which have negative values of $Q$. Visualizing regions of positive $Q$ through time will highlight the dynamics of the primary vortex, which is rotation-dominated, while leaving out the boundary layers. Later, $|\omega|$ visualizations will be presented to emphasize the dynamics of the wire’s boundary layer and secondary vortices and its interaction with the primary vortex.
@@ III.1 The distinction between weak and strong vortex regimes
Figure 2: Volume rendering visualization of the time evolution of $Q$ (Eq. 3) for a simulation in the strong vortex regime with $Re_{\Gamma}=1000$, $I_{P}=0.05$. Regions in blue indicate regions of strongly positive $Q$. The wire is shown as a solid black object.
Figure 3 depicts a similar volume rendering of $Q$ in the weak vortex regime ($Re_{\Gamma}=1000$, $I_{P}=0.25$). In this regime, the primary vortex does not “pre-sense” the wire, instead interacting with it solely when it is in close proximity to the vortex core. The vortex swiftly traverses the tube, causing significant deformation to the core. This deformation generates upwards and downwards-propagating waves. As time progresses, the wire completes its passage through the primary vortex tube, leaving behind a minor wake and a distorted tube. Surprisingly, despite the distortion, the “weak” vortex tube displays better resilience compared to the scenario involving a strong vortex. By the end of the BVI process, the waves have transversed the tube and a distorted vortex tube remains behind. Counter-intuitively this means that at $Re_{\Gamma}=1000$, a vortex tube endures the “cut” of its core more effectively when encountering a weaker cutting object. Primarily, this outcome emerges because the interaction with a slower wire is prolonged, allowing the vortex more time to engage with secondary vorticity originating from the wire. This extended interaction disrupts the primary vortex to a greater extent than the short interaction of the weak vortex regime.
Figure 3: Volume rendering visualization of the time evolution of $Q$ (Eq. 3) for a simulation in the weak vortex regime with $Re_{\Gamma}=1000$, $I_{P}=0.25$. Regions in blue indicate regions of strongly positive $Q$. The wire is shown as a solid black object.
@@
Figure 4: Volume rendering visualization of $Q$ (Eq. 3) after the wire has passed through the cylinder axis for $Re_{\Gamma}=1000$, and $I_{P}=0.05$, $0.1$, $0.15$ and $0.25$ (left to right). Regions in blue indicate regions of strongly positive $Q$. The wire is shown as a solid black object.
III.2 The strong vortex regime ($I_{P}=0.05)$
As the weak and strong vortex regimes show very different dynamics, they will be discussed separately, taking the lowest and highest values of $I_{P}$ as paradigmatic for each regime. This subsection discusses the strong vortex regime, with $I_{P}=0.05$.
@@
To investigate the role of secondary vorticity in the vortex-wire interaction, Figure 5 presents volumetric representations of the vorticity magnitude $|\omega|$ at the early stages of interaction for both $Re_{\Gamma}=1000$ and $Re_{\Gamma}=3000$. This stage of the interaction is rich in details, and the visualizations align with the experimental images and descriptions in Ref. [17]: as the vortex approaches the tube, vorticity from its boundary layer detaches, and begins to curl around the primary tube. This is marked with a blue arrow at $t\Gamma/D^{2}=740$. Even before the wire approaches the vortex core, this detached vorticity loops once around the main vortex, distorting the core, as seen at $t\Gamma/D^{2}=940$. This is because the wrapped vorticity induces velocities which commence a compression or “necking” of the core. This phenomena is accentuated as the cylinder approaches the primary vortex. In the impact zone, the vortex and cylinder have oppositely signed vorticity, which, when pressed together further thins the neck region. The primary vorticity increases to conserve circulation, and a region with increased vorticity magnitude becomes apparent at $t\Gamma/D^{2}=1140$ (marked with a black arrow) and $t\Gamma/D^{2}=1340$. At the same time, the detached vorticity propagates further along the vortex axis, nearly reaching the computational domain’s periodic boundaries as seen in the last snapshot at $t\Gamma/D^{2}=1540$.
@@
Figure 5: Volume rendering visualization of the time evolution of $|\omega|$ during the early stages for a simulation in the strong vortex regime $I_{P}=0.05$ with $Re_{\Gamma}=1000$ (top row) and $Re_{\Gamma}=3000$ (bottom row).
@@
The observations detailed here mirror those documented experimentally in Ref. [17], while expanding upon the experimental groundwork in several crucial ways. Firstly, Figure 5 portrays vorticity three-dimensionally rather than photographs of dye visualization. This representation enables us to illustrate the propagation of secondary vorticity through the axis, a phenomenon which is not clear in the experiments, as well as showing the new vortex filaments produced. Secondly, we demonstrate that secondary vorticity arises without the need to introduce a source of turbulence, either through background flow or vibrations of the structure. Even starting from simple initial conditions, such as a Lamb-Oseen vortex, the secondary vorticity originating solely from the wire’s boundary layer is enough to faithfully reproduce the significant disruptions in the primary vortex seen experimentally. Lastly, we note that our results are roughly symmetrical around the cutting plane in our results due to the absence of axial flow. This shows that the presence of a weak axial flow does not change the phenomenology of strong vortex BVI very much. We also note that for both Reynolds numbers, the flow phenomenology at the early stages is very similar, including the extent of wrapping and speed of propagation of secondary vorticity around the primary vortex, even if the vorticity itself appears much sharper due to the reduced viscosity. This facilitates the comparison to the experiment.
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Figure 6: Vorticity magnitude $|\omega|$ at the $z=0$ plane evolution just before impact in the strong vortex regime with $I_{P}=0.05$ at $Re_{\Gamma}=1000$ (top row) and $Re_{\Gamma}=3000$ (bottom row). Snapshots are at $t\Gamma/D^{2}=940$ (left), $1140$ (middle) and $1340$ (right). Abbreviations: “DSV” detached secondary vorticity, “PVN” primary vortex necking.
To focus on the early stages of the collision, Figure 6 provides more qualitative data by representing the vorticity magnitude plotted at the central $x$-$y$ plane at three time instants as the wire is approaching the tube. The early detachment of secondary vorticity is seen in the left panels of Figure 6. Significant detachment and wrapping are present even if the wire is several diameters away from the core. The detached vorticity appears much sharper at $Re_{\Gamma}=3000$, but the degree of wrapping is similar, as will be confirmed below. The middle panels show how the primary vortex has already been distorted before the wire begins to penetrate into the vortex. The wrapped vorticity induces velocities which begin to thin out the center of the vortex. This process is accentuated as the physical interaction begins- the core is further compressed producing a “necking” seen in the right panels. We note that at $Re_{\Gamma}=1000$ even if the core is compressed, the vorticity intensification is smaller (or non-present). The reasons for this will be analyzed below.
Another view of the process is shown in Figure 7 which presents the vorticity magnitude plotted at the central $x$-$z$ plane. The snapshots in the left column confirm that the dynamics of the secondary vorticity detachment process are approximately Reynolds independent: the degree of wrapping is the same for both Reynolds number. The middle panel also shows how the vortex is distorted by the wrapped vorticity before the close interaction begins. Finally, the right panel shows the degree of “necking” present in the primary vortex as the body comes close to the vortex. The middle and right panels also reveal the main effect of $Re_{\Gamma}$: while at $Re_{\Gamma}=1000$ the primary vortex remains more-or-less smooth, at $Re_{\Gamma}=3000$ the vorticity profile is highly deformed as vorticity concentrates in thin sheets outside the center. We also note that the core has been substantially displaced from its original position.
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Figure 7: Vorticity magnitude $|\omega|$ at the $y=0$ mid-plane evolution for a simulation in the strong vortex regime at $I_{P}=0.05$ with $Re_{\Gamma}=1000$ (top row) and with $Re_{\Gamma}=3000$ (bottom row). Snapshots are at $t\Gamma/D^{2}=740$ (left), $940$ (middle), $1340$ (right). Abbreviations: “DSV” detached secondary vorticity, “PVN” primary vortex necking.
However, these differences do not result in distinct outcomes of the BVI. Figure 8 shows the vorticity modulus for the late stages of the interaction, once the wire has cut the vortex core. As discussed in Section III.1, in the strong vortex case the vortex remnants are much weaker than the original primary vortex. Instead, the flow appears to consist of a turbulent cloud with vortices oriented in many directions.
@@
Figure 8: Volume visualization of the time evolution of $|\omega|$ during the late stages for a simulation in the strong vortex regime $I_{P}=0.05$ with $Re_{\Gamma}=1000$ (top row) and $Re_{\Gamma}=3000$ (bottom row).
Another view of the remnants is shown in Figure 9, which presents the vorticity at the $y=0$ plane at a late stage. For $Re_{\Gamma}=1000$, the remnants of the primary vortex can still be distinguished. It has been displaced, six to seven vortex radii from its initial position, and has a circulation which has been reduced to a quarter of its original strength. For $Re_{\Gamma}=3000$, the remnants are extremely weak, and the circulation cannot be adequately measured as a single remnant cannot be identified. In both cases, the remnants will be dissipated by viscosity.
Refer to captionRefer to caption
Figure 9: Vorticity magnitude $|\omega|$ at the $y=0$ mid-plane evolution and $t\Gamma/D^{2}=2140$ for a simulation in the strong vortex regime at $I_{P}=0.05$ with $Re_{\Gamma}=1000$ (left) and with $Re_{\Gamma}=3000$ (right). Abbreviations: “PVR” primary vortex remnants.
III.3 The weak vortex regime ($I_{P}=0.25$)